The total potential energy describing the NaCl lattice is given by: Utot = [ ( -(αe^2) / (4πε 0r) ) + ( A / r^8 ) ], where a is the Madelung constant (a=1.75 for NaCl) and A is a positive constant.

1. What is the physical meaning of the two terms of Utot? Let r0 be the value of the ionic separation r for which Utot is minimum.

2. Deduce an expression of A as a function of r0. Using r0=0.281nm, extract a value of Utot for r=r0.

The quantity –Utot is the energy required to remove a Na+ ion and a Cl- ion from the lattice. To form a pair of neutral atoms from this pair of ions involves a release of 5.14eV (the ionization energy of Na) and the expenditure of 3.61eV (the electron affinity of Cl).

3. What is the energy required to remove a pair of neutral Na and Cl atoms from the crystal? The experimental value for this quantity is 6.39eV. How well does your calculation agree?
Answer #1

2 O given ( ) ( 0 L + 00 = - 1 2 4T to no I it represent attractive potential it represent the repusive potential present Ten2 do? 4TT to По +II со 32Тео put value of A in ear (1 - Usotay = - Qe2 + 4-то по de este 32 тео та put ит = axo1 ) - xo2 4TT- 389 - - (7x1.75% 16X16 JX9 X 10 X 10 8x 0.281 8109 - 20 Uttel = - 125.55 x10° dvided by 1:6 x 09. Uporal = -7.84 R ev as gdear I solved each part clearly......if u have any doubt PlZ comment.

Plz like my answer.......Thank you.